OpenStudy (anonymous):

question for calculus: A rock is dropped straight off a bridge that is 50 meters above the ground. Another person is 7 meters away on the same bridge. At what rate is the distance between the rock and the second person increasing just as the rock hits the ground?

OpenStudy (jamesj):

Let the vertical position of the rock be s(t). Then the distance from the other person by Pythagorus' theorem is \[ f(t) = \sqrt{7^2 + s(t)^2} \] Now, what is the function s(t)?

OpenStudy (jamesj):

Given the way we're measuring distance s(0) = 0. What is s(t) in general? At what rate is the rock accelerating? Hence what is its velocity and hence what is its position?

OpenStudy (anonymous):

would the function s(t)= 0 as well?

OpenStudy (jamesj):

clearly not because the rock changes position.

OpenStudy (anonymous):

I think the rate of acceleration is -9.8 m/sec2 but I could be wrong

OpenStudy (jamesj):

yes, call that g. So at time t, where is the rock? How far has it fallen?

OpenStudy (anonymous):

so, at t(0), has the rock fallen yet, or is it still at zero?

OpenStudy (jamesj):

At t = 0 is the exact moment the rock is released. At the exact moment, it has not travelled any distance from the person who just released it. But there is now a force acting on the rock--gravity--and it is accelerating at a rate of g. Let's break it down then. Given that the acceleration is constant, what is the rock's speed at time t, given that its velocity at time t=0 is also zero.

OpenStudy (anonymous):

so, the equation for velocity would be v(t)=-9.8t+c, but since c=0, would v(t) just equal -9.8t?

OpenStudy (jamesj):

yes. Now, what then is s(t)?

OpenStudy (anonymous):

s(t) would equal \[-9.8t ^{2} /2 +c\] ?

OpenStudy (jamesj):

Yes and c equals what in this case?

OpenStudy (anonymous):

would that equal 50, since that's the distance is travelled?

OpenStudy (jamesj):

No. At t = 0, s(t) = 0. i.e., s(0) = 0. Hence c = ...

OpenStudy (anonymous):

0?

OpenStudy (jamesj):

yes. We're measuring vertical distance from the bridge. That is why the function f(t) that shows the distance the rock has travelled has the form it has. Draw a diagram and make sure you understand where f(t) comes from.

OpenStudy (jamesj):

So what you now have is that the distance of the rock from the other person is \[ f(t) = \sqrt{7^2 + s(t)^2} = \sqrt{49 + (9.8^2/2^2)t^4} = \sqrt{49 + at^4} \] where a = 9.8^2/2^2 = 24.01

OpenStudy (jamesj):

The question now asks what's the rate of change of this function when the rock hits the ground. To answer now you need to 1. Find the time T when s(t=T) = -50, i.e., the ground. 2. Find the derivative of f, f'(t) 3. Evaluate the derivative at t=T, f'(T)

OpenStudy (jamesj):

I'll leave you to it.

OpenStudy (anonymous):

okay. For 1, we would set the s(t)=-50, and when we figure it out, I think the answer is closest to t=3.19

OpenStudy (anonymous):

as for f'(t), I think it equals \[2\sqrt{49+at ^{4)}}^{1/2} \times 4t ^{3} \] ?

OpenStudy (anonymous):

and the last part, i'm not too sure about

OpenStudy (jamesj):

For the derivative, use the chain rule. You haven't got it right.

OpenStudy (anonymous):

\[1/2 \sqrt{49+at ^{4}}^{-1/2} \times 4t ^{3}\]

OpenStudy (jamesj):

You dropped an "a"

OpenStudy (jamesj):

\[ f'(t) = \frac{4at^3}{2\sqrt{49+at^4}} \]

OpenStudy (anonymous):

oh okay. Now, would this function be used as dz/dt in the equation?

OpenStudy (jamesj):

What is z? Ask yourself, what is f(t)?

OpenStudy (anonymous):

z would be the original f(t) \[\sqrt{49+at ^{4}}\]

OpenStudy (jamesj):

If you insist. In which case, if z = f(t), then clearly dz/dt = df/dt = f'(t)

OpenStudy (anonymous):

Okay, so then when we plug that back into the original equation, we're left to solve for dy/dt ?

OpenStudy (jamesj):

What is it we're trying to measure, i.e., what is the question asking? "At what rate is the distance between the rock and the second person increasing just as the rock hits the ground?" Now, as f(t) is the distance between the rock and the second person, we want to know f'(t) when the rock hits the ground.

OpenStudy (jamesj):

So you want to know the value of f'(t) for the appropriate value of t.

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

Yes, now it does

OpenStudy (anonymous):

Would we use the same value of for t what we found before, 3.19?

OpenStudy (jamesj):

for what value of t does the rock hit the ground?

OpenStudy (anonymous):

I think 3.19?

OpenStudy (jamesj):

How did you derive t = 3.19?

OpenStudy (anonymous):

well, when we figured out s(t)=-50, the answer I got was approximately equal to 3.19

OpenStudy (jamesj):

Right. So that must be the right value of t.

OpenStudy (anonymous):

okay, so then we plug that into the f'(t) equation and solve for it?

OpenStudy (jamesj):

solve it, no. But evaluate it, yes.

OpenStudy (jamesj):

The final answer to the problem is the value of f'(3.19)

OpenStudy (jamesj):

So do yourself a favor now. Go back to the beginning of the problem, and a blank piece of paper, and write out all of the steps yourself and make sure you understand where they come from. When you can do that without looking at this page on OpenStudy or your earlier versions, then you really understand that problem. I strongly recommend doing that. You'll really learn that way.

OpenStudy (anonymous):

Okay, thank you for your help